Circular No. 8509

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005bh IN UGC 6495 Further to IAUC 8506, T. Puckett and L. Cox report the discovery

of an apparent supernova (mag 17.3) on an unfiltered CCD frame taken with the 0.60-m automated supernova patrol telescope on Apr. 10.10 UT. The new object was confirmed (also at mag 17.3) on CCD frames taken by M. BenDaniel, Observatorio del Teide, Canary Islands, with the ‘SLOOH’ 0.35-m reflector on Apr. 10.91. SN 2005bh is located at α = 11^{h}29^{m}44^{s}.00, δ = +22^{o}07′39′′.1 (equinox 2000.0), which is 16′′.7 west and 3′′.7 north of the center of UGC 6495. Nothing is present at this location on images taken by Puckett on 2002 Feb. 13 and 2003 Mar. 15 (limiting mag ∼ 20.0).

SUPERNOVAE 2005bb AND 2005bf N. Morrell, M. Hamuy, G. Folatelli, and C. Contreras, Carnegie Super-

nova Project, report that spectroscopic observations (range 380–930 nm) of SN 2005bf (cf. IAUC 8507) and SN 2005bb (cf. IAUC 8504) were obtained on Apr. 7.08 and 7.23 UT, respectively, with the Las Campanas 2.5-m du Pont telescope (+ WFCCD spectrograph). SN 2005bf is a type-Ic supernova a few days before maximum light, with very strong Si II 635.5-nm and a weak He I 587.6-nm P Cyg line, among other features. SN 2005bb is of type II, highly reddened, with strong P-Cyg profiles in Hα and Ca II 857.9-nm, and weak He I 587.6-nm. The expansion velocity, derived from the minimum of the Hα absorption, is 6000 km/s, assuming for the host galaxy the recession velocity given in the NED database.

M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithsonian Center for Astrophysics; and T. Matheson, National Optical Astronomy Observatory, report that a spectrum (range 340–730 nm) of SN 2005bf, obtained on Apr. 9.23 UT by H. Landt with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a young type-Ic supernova with a blue continuum and relatively shallow absorption lines. The spectrum is similar to early spectra of SN 1987M (Filippenko et al. 1990, A.J. 100, 1575).

V378 SERPENTIS Visual magnitude estimates: Apr. 4.148 UT, 11.7 (P. Schmeer, Bis-

chmisheim, Germany); 5.741, 11.6 (S. Yoshida, Ibaraki, Japan).

2005 April 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green